At May 31, 2020, the accounts of Lopez Company show the following.

1. May 1 inventories—finished goods $14,800, work in process $17,600, and raw materials $8,600.
2. May 31 inventories—finished goods $9,600, work in process $17,000, and raw materials $8,000.
3. Debit postings to work in process were direct materials $64,300, direct labor $51,400, and manufacturing overhead applied $42,100. (Assume that overhead applied was equal to overhead incurred.)
4.

Sales revenue totaled $217,000.

Prepare an income statement for May through gross profit.
Answer #1

Working notes:

To arrive grss profit first we need to calculate COGS

COGS = Opening finished goods inventory+COGM-Closing finishd goods inventory

So we have to calculate COGM.

COGM = Opening work in progress+Total manufacturing cost(TMC)-Closing work in progress

We know work in progress fgures in the question but we have to calculate Total Manufacturing cost (TMC)

Total Manufacturing cost (TMC) = Direct raw matearial used+Direct labour+Manufacturing overhead.

Direct raw matearial used = Opening raw material+Purchases-Closing raw matearial

Let us arrive the amounts

Direct raw matearial used = 64,900 (8600+64,300-8000)

Total Manufacturing cost (TMC) =120,400 (64,900+51,400+42,100)

COGM = 121,000 (17,600+120,400-17,000)

COGS = 126,200 (14,800+121,000-9,600)

Income Statement of Lopez company for the month of May 2020 is as below.

Sales Less: COGS Gross profit 2,17,000 1,26,200 90,800